Sally is 3 years older than Helen. The product of their ages is 70. How old is Helen?

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

Sally is 3 years older than Helen.
s = h + 3
The product of their ages is 70.
s * h = 70
How old is Helen?
in the above equation, replace s with (h+3)
(h+3)*h = 70
h^2 + 3h = 70
our old friend, the quadratic equation!!
h^2 + 3h - 70 = 0
You can use the quadratic equation; a=1; b=3; c=-70, but this will factor to
(h+10)(h-7) = 0
I'm sure Helen is a positive person, therefore
h = 7 yrs
:
:
Obviously Sally will be 10, it checks out; 10 * 7 = 70

Answer by Alan3354(69443)

Sally is 3 years older than Helen. The product of their ages is 70. How old is Helen?
==================
Find a pair of factors of 70 that differ by 3.
---
Or, make a quadratic, then find a pair of factors of 70 that differ by 3 to factor it.